OpenStudy (anonymous):

What is the value of s in the equation 3r = 10 + 5s, when r = 10? 4 8 100 200

OpenStudy (zpupster):

3(10) = 10 + 5s plug in 10 30 - 10 = 10-10 + 5s add 10 to both sides 20/5 = 5s/5 ??=s